OpenStudy (anonymous):

help someone, please?! i dont want the answer, i want guidance. The expression dy/dx = x(3√y) gives the slope at any point on the graph of the function f(x) where f(2) = 8.

OpenStudy (anonymous):

A. Write the equation of the tangent line to f(x) at point (2, 8). what am i supposed to do? differentiate?

OpenStudy (anonymous):

@mathmale could you please help me?

OpenStudy (anonymous):

the equation came out wrong. it's supposed to look like this: \[dy/dx=x(\sqrt[3]{y})\]

OpenStudy (anonymous):

what are the initial steps i should take?

OpenStudy (anonymous):

@AccessDenied are you going to help me?

OpenStudy (accessdenied):

One note: y = f(x). Because dy/dx = the slope of f(x) at a point. We want to find the equation for a tangent line. So, we want two pieces of information: The point, which is given: (2, 8) The slope at that point, which we use dy/dx = x cuberoot(y) and our point to find.

OpenStudy (anonymous):

@agent0smith help please?

OpenStudy (accessdenied):

We can plug in our point information into the formula for dy/dx, which will give us the slope at the point. Then we use point-slope formula to get the equation of the line: y - y0 = m(x - x0) (x0, y0) = (2, 8). m = dy/dx(2,8)

OpenStudy (anonymous):

so do we just plug in the point into the equation then make an equation?

OpenStudy (anonymous):

i got it! i think... would the equation of the tangent line be y=4x?

OpenStudy (accessdenied):

That looks good! :)

OpenStudy (accessdenied):

Sorry the site had a moment and I went to do some other things in the mean time; I missed the updates!

OpenStudy (anonymous):

it's okay. i solved it! thanks a million!

OpenStudy (accessdenied):

Glad to help!